# Simple fits for the neutrino luminosities from protoneutron star cooling

### v2

@@ Appendix D Tables for parameter values of the correlations between $\tau$ and $\alpha$
We report in Tab. 9 the best-fit values and the $1\,\sigma$ errors for the parameters $A$ and $B$ of the linear functions used for describing the correlations between $\tau$ (in seconds) and $\alpha$ [see Eq. (5)]:
@@
As shown in Tab. 9, the fit works better for neutrinos than for antineutrinos. Indeed, for $\bar{\nu}_{e}$ the error on the parameter $A$ is larger than its best-fit value (thus, $A$ is compatible with zero) for all the NS masses, and the same is true for $\bar{\nu}_{\mu}$ for the largest NS mass. Additionally, the quality of the fit is similar for $\nu_{\mu}$ and $\nu_{\tau}$, while it is slightly better for $\bar{\nu}_{\tau}$ compared to $\bar{\nu}_{\mu}$, since the relative error of the best-fit parameters for $\bar{\nu}_{\tau}$ is smaller. This reveals a small difference between the non-electron flavors. As a common trend, the parameter $B$ increases with the NS mass for all of the neutrino species.
@@ Appendix E Further details on the impact of convection and muons
Here we give further details on the impact of convection and muons. Since models 1.62-SFHx-c and 1.61-LS220-c stop before the product $tL_{\nu_{i}}$ for all neutrino species is reduced to a value of 0.15 of the maximum, to make the comparison on a solid ground in this Section we consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). In this way, we take into account results for neutrinos and antineutrinos of all flavors up to the time when they reach the same reduction factor. Moreover, since non-electron flavors show in general a similar behavior and in simulations without muons, $\tau$ and $\mu$ neutrinos behave exactly in the same way, for all the simulations considered in this Section we report values related only to $\nu_{e}$, $\bar{\nu}_{e}$, $\nu_{\mu}$ and $\bar{\nu}_{\mu}$.
@@
We list the best-fit parameter values and their $1\,\sigma$ errors for the $1.62~{}M_{\odot}$ models with different EoS and different ingredients of the input physics for the luminosities of neutrinos in Tab. 10 and of antineutrinos in Tab. 11. A better visualization of the change in the best-fit parameters can be obtained by plotting them as a function of the PNS mass for the different EoSs. As an example, we show in Fig. 16 the best-fit parameters for $\bar{\nu}_{e}$ as a function of $M_{\rm NS}$, for DD2 (left panels) and SFHo (right panels). Here, black dots are the values of the best-fit parameters obtained from simulations including both convection and muons, red dots are related to simulations without convection and blue dots to simulations without muons.
@@
Table 10 and Tab. 11, as well as Fig. 16, show that, as a general trend, in simulations without convection $\tau$ becomes larger and $n$ smaller, i.e., the luminosity suppression starts at later times and it is slower. The only exception is found for $\tau_{\bar{\nu}_{e}}$ in model 1.62-DD2-c, which is smaller than $\tau_{\bar{\nu}_{e}}$ in 1.62-DD2 (see the red dot in the third panel panel from top on the left of Fig. 16). This behavior is confirmed by inspecting the upper left panel in Fig. 13, where the orange line (without convection) is peaked at earlier times compared to the blue (benchmark case) and the green (without muons) lines. As shown by the upper left panel in Fig. 14, even if $\tau_{\bar{\nu}_{e}}$ is smaller than in the benchmark case, the fit well reproduces the data, because the interplay between $\tau_{\bar{\nu}_{e}}$ and a much smaller $n_{\bar{\nu}_{e}}$ (compared to the complete case) well describes the slightly longer cooling time. Therefore, for $\bar{\nu}_{e}$ we observe a mathematical peculiarity connected to the fit function, reacting to the fact that $tL_{\bar{\nu}_{e}}$ is peaked at earlier times in the absence of convection (see the upper left panel in Fig. 13), i.e. smaller $\tau_{\bar{\nu}_{e}}$, and it is characterized by a milder exponential suppression, i.e. smaller $n$, leading to a longer cooling time. On the other hand, in the absence of convection $\alpha$ becomes smaller in the case of DD2 (see the red dot in the second panel from top on the left of Fig. 16) and larger in the case of SFHo (see the red dot in the second panel from top on the right of Fig. 16), describing a change in the power-law behavior in the early cooling phase. In all the cases, the best-fit parameters in the absence of convection lie well outside the $1\,\sigma$ confidence band found for benchmark simulations, stressing the strong impact of convection on the neutrino signal.
@@
The weaker impact of muons on the neutrino signal is highlighted by the small changes in the best-fit parameters obtained from simulations without muons. In this case, for both DD2 and SFHo, in simulations without muons $\tau$ becomes slightly lower and $n$ is approximately equal or slightly larger for neutrinos and antineutrinos of all flavors (compare the first with the third line in the first two data blocks of Tab. 10 and Tab. 11, as well as the black and blue dots in Fig. 16). This means that in simulations without muons the suppression in the luminosity starts slightly before and it is a bit faster than in the full-physics cases. In contrast, $\alpha$ tends to increase for DD2 and to decrease for SFHo, even if the change in all cases is much smaller compared to the changes induced by the absence of convection.
@@
Since simulations without muons are not available for SFHx and LS220 and for them $X_{\nu_{i}}^{\rm fin}>0.15$ in the absence of convection, we do not show the best-fit parameter values as a function of the NS mass in these two cases, but we only summarize the values of the best-fit parameters and their errors in the last two data blocks of Tab. 10 (for the electron flavor) and Tab. 11 (for the muon flavor), obtained by considering simulation data up to $t_{\nu_{i},{\rm Co}}$. Even if the nominal values of the best-fit parameters in the benchmark simulations slightly change when switching from $t_{\nu_{i},{\rm c}}$ to $t_{\nu_{i},{\rm Co}}$, the impact of convection on simulations with SFHx and SFHo is similar, with an increase in $\tau$, and a decrease in $C$ and in $n$ in absence of convection. As expected, convection strongly affects also simulations with LS220. In this case, given the worse quality of the fit, neglecting convection leads to completely different values of the best-fit parameters compared to the benchmark case. Indeed, simulations without convection show positive values of $\alpha$, much larger values of $\tau$ and $n$, and drastically reduced values of $C$ compared to the complete-physics case, with all of these parameter values more closely related to the true magnitude and exponential decay time of the neutrino luminosities.
@@
Table 10: Best-fit parameter values for $L_{\nu_{i}}=C\,t^{-\alpha}\,e^{-(t/\tau)^{n}}$ for $\nu_{e}$ (left) and $\bar{\nu}_{e}$ (right) for $M_{\rm NS}=1.62~{}M_{\odot}$ and different EoS, considering both convection and muons (upper lines), without convection (labelled with the suffix -c) and without muons (labelled with the suffix -m). We consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). Simulations with SFHx and LS220 without muons are not available.
@@
Table 11: Best-fit parameter values for $L_{\nu_{i}}=C\,t^{-\alpha}\,e^{-(t/\tau)^{n}}$ for $\nu_{\mu}$ (left) and $\bar{\nu}_{\mu}$ (right) for $M_{\rm NS}=1.62~{}M_{\odot}$ and different EoS, considering both convection and muons (upper lines), without convection (labelled with the suffix -c) and without muons (labelled with the suffix -m). We consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). Simulations with SFHx and LS220 without muons are not available.
Refer to captionRefer to caption

### v3

@@ Appendix D Tables for parameter values of the correlations between $\tau$ and $\alpha$
We report in Table 9 the best-fit values and the $1\,\sigma$ errors for the parameters $A$ and $B$ of the linear functions used for describing the correlations between $\tau$ (in seconds) and $\alpha$ [see Eq. (5)]:
@@
As shown in Table 9, the fit works better for neutrinos than for antineutrinos. Indeed, for $\bar{\nu}_{e}$ the error on the parameter $A$ is larger than its best-fit value (thus, $A$ is compatible with zero) for all the NS masses, and the same is true for $\bar{\nu}_{\mu}$ for the largest NS mass. Additionally, the quality of the fit is similar for $\nu_{\mu}$ and $\nu_{\tau}$, while it is slightly better for $\bar{\nu}_{\tau}$ compared to $\bar{\nu}_{\mu}$, since the relative error of the best-fit parameters for $\bar{\nu}_{\tau}$ is smaller. This reveals a small difference between the non-electron flavors. As a common trend, the parameter $B$ increases with the NS mass for all of the neutrino species.
@@ Appendix E Further details on the impact of convection and muons
Here we give further details on the impact of convection and muons. Since models 1.62-SFHx-c and 1.61-LS220-c stop before the product $tL_{\nu_{i}}$ for all neutrino species is reduced to a value of 0.15 of the maximum, to make the comparison on a solid ground in this appendix we consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). In this way, we take into account results for neutrinos and antineutrinos of all flavors up to the time when they reach the same reduction factor. Moreover, since non-electron flavors show, in general, a similar behavior and in simulations without muons, $\tau$ and $\mu$ neutrinos behave exactly in the same way, for all the simulations considered in this Section we report values related only to $\nu_{e}$, $\bar{\nu}_{e}$, $\nu_{\mu}$ and $\bar{\nu}_{\mu}$.
@@
We list the best-fit parameter values and their $1\,\sigma$ errors for the $1.62~{}M_{\odot}$ models with different EoS and different ingredients of the input physics for the luminosities of neutrinos in Table 10 and of antineutrinos in Table 11. A better visualization of the change in the best-fit parameters can be obtained by plotting them as a function of the PNS mass for the different EoSs. As an example, we show in Fig. 16 the best-fit parameters for $\bar{\nu}_{e}$ as a function of $M_{\rm NS}$, for DD2 (left panels) and SFHo (right panels). Here, black dots are the values of the best-fit parameters obtained from simulations including both convection and muons, red dots are related to simulations without convection and blue dots to simulations without muons.
@@
Tables 10 and 11, as well as Fig. 16, show that, as a general trend, in simulations without convection $\tau$ becomes larger and $n$ smaller, i.e., the luminosity suppression starts at later times and it is slower. The only exception is found for $\tau_{\bar{\nu}_{e}}$ in model 1.62-DD2-c, which is smaller than $\tau_{\bar{\nu}_{e}}$ in 1.62-DD2 (see the red dot in the third panel from top on the left of Fig. 16). This behavior is confirmed by inspecting the upper left panel in Fig. 13, where the orange line (without convection) is peaked at earlier times compared to the blue (benchmark case) and the green (without muons) lines. As shown by the upper left panel in Fig. 14, even if $\tau_{\bar{\nu}_{e}}$ is smaller than in the benchmark case, the fit well reproduces the data, because the interplay between $\tau_{\bar{\nu}_{e}}$ and a much smaller $n_{\bar{\nu}_{e}}$ (compared to the complete case) well describes the slightly longer cooling time. Therefore, for $\bar{\nu}_{e}$ we observe a mathematical peculiarity connected to the fit function, reacting to the fact that $tL_{\bar{\nu}_{e}}$ is peaked at earlier times in the absence of convection (see the upper left panel in Fig. 13), i.e. smaller $\tau_{\bar{\nu}_{e}}$, and it is characterized by a milder exponential suppression, i.e. smaller $n$, leading to a longer cooling time. On the other hand, in the absence of convection $\alpha$ becomes smaller in the case of DD2 (see the red dot in the second panel from top on the left of Fig. 16) and larger in the case of SFHo (see the red dot in the second panel from top on the right of Fig. 16), describing a change in the power-law behavior in the early cooling phase. In all the cases, the best-fit parameters in the absence of convection lie well outside the $1\,\sigma$ confidence band found for benchmark simulations, stressing the strong impact of convection on the neutrino signal.
@@
The weaker impact of muons on the neutrino signal is highlighted by the small changes in the best-fit parameters obtained from simulations without muons. In this case, for both DD2 and SFHo, in simulations without muons $\tau$ becomes slightly lower and $n$ is approximately equal or slightly larger for neutrinos and antineutrinos of all flavors (compare the first with the third line in the first two data blocks of Table 10 and Table 11, as well as the black and blue dots in Fig. 16). This means that in simulations without muons the suppression in the luminosity starts slightly before and it is a bit faster than in the full-physics cases. In contrast, $\alpha$ tends to increase for DD2 and to decrease for SFHo, even if the change in all cases is much smaller compared to the changes induced by the absence of convection.
@@
Since simulations without muons are not available for SFHx and LS220 and for them $X_{\nu_{i}}^{\rm fin}>0.15$ in the absence of convection, we do not show the best-fit parameter values as a function of the NS mass in these two cases, but we only summarize the values of the best-fit parameters and their errors in the last two data blocks of Table 10 (for the electron flavor) and Table 11 (for the muon flavor), obtained by considering simulation data up to $t_{\nu_{i},{\rm Co}}$. Even if the nominal values of the best-fit parameters in the benchmark simulations slightly change when switching from $t_{\nu_{i},{\rm c}}$ to $t_{\nu_{i},{\rm Co}}$, the impact of convection on simulations with SFHx and SFHo is similar, with an increase in $\tau$, and a decrease in $C$ and in $n$ in absence of convection. As expected, convection strongly affects also simulations with LS220. In this case, given the worse quality of the fit, neglecting convection leads to completely different values of the best-fit parameters compared to the benchmark case. Indeed, simulations without convection show positive values of $\alpha$, much larger values of $\tau$ and $n$, and drastically reduced values of $C$ compared to the complete-physics case, with all of these parameter values more closely related to the true magnitude and exponential decay time of the neutrino luminosities.
@@
Table 10: Best-fit parameter values for $L_{\nu_{i}}=C\,t^{-\alpha}\,e^{-(t/\tau)^{n}}$ for $\nu_{e}$ (left) and $\bar{\nu}_{e}$ (right) for $M_{\rm NS}=1.62~{}M_{\odot}$ and different EoS, considering both convection and muons (upper lines), without convection (labeled with the suffix “-c”) and without muons (labeled with the suffix “-m”). We consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). Simulations with SFHx and LS220 without muons are not available.
@@
Table 11: Best-fit parameter values for $L_{\nu_{i}}=C\,t^{-\alpha}\,e^{-(t/\tau)^{n}}$ for $\nu_{\mu}$ (left) and $\bar{\nu}_{\mu}$ (right) for $M_{\rm NS}=1.62~{}M_{\odot}$ and different EoS, considering both convection and muons (upper lines), without convection (labeled with the suffix “-c”) and without muons (labeled with the suffix “-m”). We consider data up to $t_{\nu_{i},{\rm c}}$ for simulations with DD2 and SFHo and up to $t_{\nu_{i},{\rm Co}}$ for simulations with SFHx and LS220 (see Appendix A for more details). Simulations with SFHx and LS220 without muons are not available.
@@
Appendix F Equation-of-state parameters
We report in Table 12 the parameters for the symmetry energies for the EoS cases used in our work. With the customary definitions of $x=(n-n_{0})/(3n_{0})$ and the asymmetry parameter $\delta=1-2Y_{e}$, the energy per nucleon can be expressed as
$E(n)=-E_{0}+\frac{1}{2}Kx^{2}+\delta^{2}(J+Lx+\frac{1}{2}K_{\mathrm{s}}x^{2})+...\,,$ (14)
with $E_{0}$ being the binding energy of symmetric matter at saturation density, $K$ the incompressibility, $J$ the symmetry energy, $L$ the slope of the symmetry energy, and $K_{\mathrm{s}}$ the curvature of the symmetry energy.
EoS $n_{0}$ $E_{0}$ $K$ $J$ $L$ $K_{\mathrm{s}}$
[fm-3] [MeV] [MeV] [MeV] [MeV] [MeV]
DD2 0.149 16.0 243 31.7 55.0 --93.2
SFHo 0.158 16.2 245 31.6 47.1 --205
SFHx 0.160 16.2 239 28.7 23.2 --40.0
LS220 0.155 16.0 220 28.6 74.0 --24.0
Table 12: Parameter values for the energy per nucleon around the nuclear saturation density $n_{0}$ according to Eq. (14) for the EoSs used in the model simulations in our work. The values in this table are taken from Table IV in [67] and from entries for the respective EoS in the CompOSE database [68].